# FSR-VLN: Fast and Slow Reasoning for Vision-Language Navigation with Hierarchical Multi-modal Scene Graph

### v2

@@ Abstract
Visual-Language Navigation (VLN) is a fundamental challenge in robotic systems, with broad applications for the deployment of embodied agents in real-world environments. Despite recent advances, existing approaches are limited in long-range spatial reasoning, often exhibiting low success rates and high inference latency, particularly in long-range navigation tasks. To address these limitations, we propose FSR-VLN, a vision-language navigation system that combines a Hierarchical Multi-modal Scene Graph (HMSG) with Fast-to-Slow Navigation Reasoning (FSR). The HMSG provides a multi-modal map representation supporting progressive retrieval, from coarse room-level localization to fine-grained goal view and object identification. Building on HMSG, FSR first performs fast matching to efficiently select candidate rooms, views, and objects, then applies VLM-driven refinement for final goal selection. We evaluated FSR-VLN across four comprehensive indoor datasets collected by humanoid robots, utilizing 87 instructions that encompass a diverse range of object categories. FSR-VLN achieves state-of-the-art (SOTA) performance in all datasets, measured by the retrieval success rate (RSR), while reducing the response time by 82% compared to VLM-based methods on tour videos by activating slow reasoning only when fast intuition fails. Furthermore, we integrate FSR-VLN with speech interaction, planning, and control modules on a Unitree-G1 humanoid robot, enabling natural language interaction and real-time navigation.
@@ I INTRODUCTION
Visual-Language Navigation (VLN) is a fundamental task in embodied AI, enabling robots to operate effectively in complex real-world environments [1]. Despite significant progress in map-free VLN research and growing efforts to equip robots with visual-language reasoning [2, 3, 4, 5], existing methods remain limited in long-range spatial cognition [6], particularly for long-range navigation. A key bottleneck is the lack of persistent long-range spatial memory, which encodes, organizes, and retrieves environmental knowledge. Such memory allows robots to capture comprehensive spatial relationships and adapt to complex indoor environments [7, 8]. Without it, robots struggle to understand and reason about long-range spaces. To address this limitation, recent spatial memory-based approaches have explored embedding 2D and 3D geometric maps with semantic features [9, 10, 11, 12], integrating dense geometric reconstructions with the pre-trained zero-shot Vision Foundation Model (VFM) such as CLIP [13], OWL-ViT [14], and LSeg [15]. These approaches support open-vocabulary object-level retrieval while maintaining high geometric fidelity. Building on geometric semantic maps, methods such as Clio [16], ConceptGraph [17, 18, 19], RoboExp [18], and OpenIN [10] leverage open-vocabulary 3D scene graphs to represent long-range environments and provide semantic interfaces for prompting large language model (LLM). Parallel work, including HOVSG [20] and IRS [21], abstracts these dense maps into hierarchical structures spanning floors, rooms, and objects, supporting efficient semantic object retrieval for long-range navigation.
@@
Although 2D and 3D semantic maps and 3D scene graphs provide geometry-consistent and hierarchical spatial memory, they rely on pre-extracted visual features[3, 22], lack direct interaction with a vision-language model (VLM), and are sensitive to geometric noise. These limitations, in turn, hinder their adaptability to diverse user instructions and complex real-world navigation [23, 24].
@@
Inspired by human navigation, where individuals reason over previously observed images or objects without detailed 3D maps, image-based topological navigation has emerged as a promising alternative [25, 26, 27, 28]. By retaining raw image data, image-based topological maps facilitate interaction with LLM and VLM, such as GPT-4o and Gemini, which excel at reasoning over both visual and textual inputs [29]. Despite their high success rate in image-level navigation, existing methods lack explicit 3D geometric structure and usually rely on video captioning, which is inefficient for reasoning over long sequences [30].
@@ II-A Object Navigation with Geometric Semantic Map
Recent works have extended geometric point cloud maps by incorporating semantic features to enable object navigation (ObjNav). Systems such as OK-Robot [31], VLMap [23], OVL-MAP [32], and BeliefMapNav [33] embed features from Vision Foundation Models (VFMs, e.g., CLIP [13], OWL-ViT [14], LSeg [15]) into voxel grids. While this approach preserves geometric consistency and supports open-vocabulary queries, it suffers from several critical limitations. Despite these advances, the performance of the systems remains limited by the inherent weaknesses of VFMs and their sensitivity to 3D reconstruction noise and odometry drift. Reliance on a single map representation further restricts integration with LLMs and VLMs, which excel at cross-modal reasoning and generalization.
Building on semantic point cloud maps, scene graph-based methods represent objects and spatial concepts as nodes, with their relations encoded as edges, providing compact and expressive models of long-range environments [34, 35, 36]. This object-centric decomposition supports higher-level reasoning for navigation and manipulation [20]. While many works employ 3D scene graphs [35, 36, 37] to efficiently model large environments, most rely on closed-set semantics. To enable open-vocabulary queries, approaches such as ConceptGraphs [38], HOVSG [20], DOVSG [39], and OpenIN [40] construct open-vocabulary 3D scene graphs, which can also interface with LLM.
@@ II-B ObjNav with Image-based Topological Graph
Several recent works represent environments as image-based topological graphs to facilitate VLN. MobilityVLA [27] introduces a vision-language action (VLA) framework that leverages long-context VLM for goal frame retrieval and uses topological graphs for waypoint planning. Uni-NaVid [41] adopts a video-based VLA approach, retrieving the most visually similar image to a query object and generating robot actions in an end-to-end manner. ReMEmbR [42] emphasizes spatio-temporal memory by storing caption embeddings with pose and time metadata for retrieval-augmented reasoning. More recently, MapGPT [43] incorporates map-guided prompting and adaptive path planning, significantly improving zero-shot generalization. Astra [44] extends image-based topological maps with landmark information, while RoboHop [45] employs a zero-shot “segment servoing” strategy to reach object subgoals. Building on RoboHop’s open-set navigation pipeline, TANGO [46] generates sub-object goals through a global path planner grounded in an object-level topological graph.
Although these methods preserve rich semantic cues and achieve strong performance in image-level reasoning, they rely solely on video captioning, which is inefficient in long video contexts and often leads to mismatches between semantic and geometric information [44]. In real-world environments, however, abundant spatial information-such as room numbers, equipment labels, and other annotations can significantly improve navigation performance [42]. To leverage this, we propose a hierarchical multi-modal scene graph representation that encodes both spatial and semantic features across floors, rooms, views, and objects. Building on these multi-modal features, we introduce a fast-to-slow reasoning mechanism to enhance navigation success rates.
@@ III Method
As illustrated in Fig. 1, we first build HMSG with real-world LiDAR–camera datasets collected by Unitree-G1. Specifically, the SOTA SLAM system FAST-LIVO2 [47] is used to extract RGBD data and poses to create an instance-level open-vocabulary map. Building on an instance-level open-vocabulary map [20], we construct the Hierarchical Multi-modal Scene Graph (HMSG), which divides the environment into four levels-floor, room, view, and object. Each node is enriched with multi-modal features, including geometric attributes, semantic information, and topological connections.
During online interaction, the speech recognition module [48] converts user speech into text, which is then parsed by an LLM to extract the corresponding query. In the FSR pipeline, the query text is first grounded to candidate regions and objects through CLIP-based similarity matching and then refined via VLM-based visual verification to ensure robust and accurate navigation. By combining the HMSG representation with fast-to-slow reasoning (FSR), FSR-VLN enables efficient and interpretable navigation, particularly in long-range and cluttered real-world environments. Further details of this mechanism are provided in Section III-C. After obtaining the target object’s coordinates, the system performs path planning and whole-body control to reach the goal.
@@ III-B Hierarchical Multi-modal Scene Graph Construction
Following HOVSG [20], we first construct an open-vocabulary map at the instance-level. Objects and views are associated with rooms based on geometric overlap in top-down maps, resulting in a scene graph enriched with both geometric and semantic features.
@@ IV-A Experiment Setup
User Instructions. To evaluate the generalization of our system across diverse instructions, we follow the experimental setup of MobilityVLA [27] and crowd-source 87 user instructions in four categories: Reasoning-Free (RF) with explicit targets, Reasoning-Required (RR) requiring implicit goal inference, Small Objects (SO), which are challenging due to limited visual footprints, and Spatial Target (ST) designed to test long-range memory using room-type information. The instructions span 23, 18, 15, and 14 object categories, respectively.
@@
Baselines. We evaluated FSR-VLN against several SOTA ObjNav methods using different map representations: CLIP-based 3D voxel maps (OK-Robot [31]), CLIP-based 3D scene graphs (HOVSG [20]), and image-based topological graphs (MobilityVLA [27]). OK-Robot performs object retrieval using CLIP-based OWL-ViT [14] features. HOVSG retrieves floors, rooms, and objects based on CLIP features, with its 3D scene graph structure supporting long-range retrieval. MobilityVLA retrieves goal images from all views; as it is not open-sourced, we implemented a version using GPT-4o, first selecting the top 50 candidate frames via CLIP similarity before inference due to GPT-4o’s context limitations. OK-Robot and MobilityVLA lack room-level information and cannot interpret Spatial Target (ST) instructions referencing a specific room. For fair comparison, we consider only the object query derived from LLM-based instruction understanding for these methods and HOVSG. Retrieval is also counted as successful if objects of the same type from other rooms are returned for ST instructions. In contrast, FSR-VLN leverages room-level spatial information, and retrieval is considered successful only when the target object is in the specified room. By integrating the HMSG map with navigation reasoning, FSR-VLN jointly predicts the goal view and object, and a retrieval is successful only if both are correctly identified.
In the HM3D-SEM dataset, as osmAG-LLM, we select HOV-SG and osmAG-LLM as our baseline because they share fundamental design principles with our approach[50].
@@ IV-B Benchmark Results
OK-Robot[10] CLIP based 3D Voxel Map 0.2s 0.609 0.609 0.609 0.609 0.609 0.609 0.632 0.632 0.632 0.632 0.632
HOVSG[20] CLIP based 3D Scene Graph 0.2s 0.517 0.517 0.573 0.586 0.596 0.596 0.770 0.816 0.828 0.828 0.828
MobilityVLA[27] Image based Topological Graph + VLM 30s 0.345 0.345 0.598 0.759 0.805 0.954 - - - - -

### v3

@@ Abstract
Visual-Language Navigation (VLN) is a fundamental challenge in robotic systems, with broad applications for the deployment of embodied agents in real-world environments. Despite recent advances, existing approaches are limited in long-range spatial reasoning, often exhibiting low success rates and high inference latency, particularly in long-range navigation tasks. To address these limitations, we propose FSR-VLN, a vision-language navigation system that combines a Hierarchical Multi-modal Scene Graph (HMSG) with Fast-to-Slow Navigation Reasoning (FSR). The HMSG provides a multi-modal map representation supporting progressive retrieval, from coarse room-level localization to fine-grained goal view and object identification. Building on HMSG, FSR first performs fast matching to efficiently select candidate rooms, views, and objects, then applies VLM-driven refinement for final goal selection. We evaluated FSR-VLN across four comprehensive indoor datasets collected by humanoid robots, utilizing 87 instructions that encompass a diverse range of object categories. FSR-VLN achieves state-of-the-art (SOTA) performance in all datasets, measured by the retrieval success rate (RSR), while reducing the response time by 82% compared to VLM-based methods on tour videos by activating slow reasoning only when fast intuition fails. Furthermore, we integrate FSR-VLN with speech interaction, planning, and control modules on a Unitree-G1 humanoid robot, enabling natural language interaction and real-time navigation. The code and video can be available at https://horizonrobotics.github.io/robotlab/fsr-vln/
@@ I INTRODUCTION
Visual-Language Navigation (VLN) is a fundamental task in embodied AI, enabling robots to operate effectively in complex real-world environments [8578485]. Despite significant progress in map-free VLN research and growing efforts to equip robots with visual-language reasoning [yokoyama2024vlfm, zhang2024visionandlanguagenavigationtodaytomorrow, long2024instructnavzeroshotgenericinstruction, zhou2023navgptexplicitreasoningvisionandlanguage], existing methods remain limited in long-range spatial cognition [ramakrishnan2024does], particularly for long-range navigation. A key bottleneck is the lack of persistent long-range spatial memory, which encodes, organizes, and retrieves environmental knowledge. Such memory allows robots to capture comprehensive spatial relationships and adapt to complex indoor environments [ruan2025reactivecognitivebraininspiredspatial, Gu_2022]. Without it, robots struggle to understand and reason about long-range spaces. To address this limitation, recent spatial memory-based approaches have explored embedding 2D and 3D geometric maps with semantic features [Peng2023OpenScene, jin2025openfusionopenvocabularyrealtimescene, conceptfusion, FM-Fusion], integrating dense geometric reconstructions with the pre-trained zero-shot Vision Foundation Model (VFM) such as CLIP [radford2021learningtransferablevisualmodels], OWL-ViT [minderer2022simpleopenvocabularyobjectdetection], and LSeg [li2022languagedriven]. These approaches support open-vocabulary object-level retrieval while maintaining high geometric fidelity. Building on geometric semantic maps, methods such as Clio [Maggio2024Clio], ConceptGraph [gu2023conceptgraphsopenvocabulary3dscene, jiang2024roboexpactionconditionedscenegraph, tang2025openinopenvocabularyinstanceorientednavigation], RoboExp [jiang2024roboexpactionconditionedscenegraph], and OpenIN [jin2025openfusionopenvocabularyrealtimescene] leverage open-vocabulary 3D scene graphs to represent long-range environments and provide semantic interfaces for prompting large language model (LLM). Parallel work, including HOVSG [werby23hovsg] and IRS [chen2025irsinstancelevel3dscene], abstracts these dense maps into hierarchical structures spanning floors, rooms, and objects, supporting efficient semantic object retrieval for long-range navigation.
@@
Although 2D and 3D semantic maps and 3D scene graphs provide geometry-consistent and hierarchical spatial memory, they rely on pre-extracted visual features[zhang2024visionandlanguagenavigationtodaytomorrow, pekkanen2025visuallanguagegridmapscapture], lack direct interaction with a vision-language model (VLM), and are sensitive to geometric noise. These limitations, in turn, hinder their adaptability to diverse user instructions and complex real-world navigation [huang23vlmaps, raychaudhuri2025semanticmappingindoorembodied].
@@
Inspired by human navigation, where individuals reason over previously observed images or objects without detailed 3D maps, image-based topological navigation has emerged as a promising alternative [savinov2018semiparametric, shah2023gnmgeneralnavigationmodel, xu2024mobility, 10610234]. By retaining raw image data, image-based topological maps facilitate interaction with LLM and VLM, such as GPT-4o and Gemini, which excel at reasoning over both visual and textual inputs [li2025survey]. Despite their high success rate in image-level navigation, existing methods lack explicit 3D geometric structure and usually rely on video captioning, which is inefficient for reasoning over long sequences [anwar2024remembrbuildingreasoninglonghorizon].
@@ II-A Object Navigation with Geometric Semantic Map
Recent works have extended geometric point cloud maps by incorporating semantic features to enable object navigation (ObjNav). Systems such as OK-Robot [Liu_2024], VLMap [huang23vlmaps], OVL-MAP [10879413], and BeliefMapNav [zhou2025beliefmapnav3dvoxelbasedbelief] embed features from Vision Foundation Models (VFMs, e.g., CLIP [radford2021learningtransferablevisualmodels], OWL-ViT [minderer2022simpleopenvocabularyobjectdetection], LSeg [li2022languagedriven]) into voxel grids. While this approach preserves geometric consistency and supports open-vocabulary queries, it suffers from several critical limitations. Despite these advances, the performance of the systems remains limited by the inherent weaknesses of VFMs and their sensitivity to 3D reconstruction noise and odometry drift. Reliance on a single map representation further restricts integration with LLMs and VLMs, which excel at cross-modal reasoning and generalization.
Building on semantic point cloud maps, scene graph-based methods represent objects and spatial concepts as nodes, with their relations encoded as edges, providing compact and expressive models of long-range environments [armeni20193d, rosinol20203d, hughes2022hydra]. This object-centric decomposition supports higher-level reasoning for navigation and manipulation [werby23hovsg]. While many works employ 3D scene graphs [rosinol20203d, hughes2022hydra, greve2024collaborative] to efficiently model large environments, most rely on closed-set semantics. To enable open-vocabulary queries, approaches such as ConceptGraphs [gu2024conceptgraphs], HOVSG [werby23hovsg], DOVSG [yan2025dynamicopenvocabulary3dscene], and OpenIN [tang2025openin] construct open-vocabulary 3D scene graphs, which can also interface with LLM.
@@ II-B ObjNav with Image-based Topological Graph
Several recent works represent environments as image-based topological graphs to facilitate VLN. MobilityVLA [xu2024mobility] introduces a vision-language action (VLA) framework that leverages long-context VLM for goal frame retrieval and uses topological graphs for waypoint planning. Uni-NaVid [zhang2024uni] adopts a video-based VLA approach, retrieving the most visually similar image to a query object and generating robot actions in an end-to-end manner. ReMEmbR [anwar2025remembr] emphasizes spatio-temporal memory by storing caption embeddings with pose and time metadata for retrieval-augmented reasoning. More recently, MapGPT [chen2024mapgpt] incorporates map-guided prompting and adaptive path planning, significantly improving zero-shot generalization. Astra [chen2025astra] extends image-based topological maps with landmark information, while RoboHop [garg2024robohop] employs a zero-shot “segment servoing” strategy to reach object subgoals. Building on RoboHop’s open-set navigation pipeline, TANGO [podgorski2025tango] generates sub-object goals through a global path planner grounded in an object-level topological graph.
Although these methods preserve rich semantic cues and achieve strong performance in image-level reasoning, they rely solely on video captioning, which is inefficient in long video contexts and often leads to mismatches between semantic and geometric information [chen2025astra]. In real-world environments, however, abundant spatial information-such as room numbers, equipment labels, and other annotations can significantly improve navigation performance [anwar2025remembr]. To leverage this, we propose a hierarchical multi-modal scene graph representation that encodes both spatial and semantic features across floors, rooms, views, and objects. Building on these multi-modal features, we introduce a fast-to-slow reasoning mechanism to enhance navigation success rates.
@@ III Method
As illustrated in Fig. 1, we first build HMSG with real-world LiDAR–camera datasets collected by Unitree-G1. Specifically, the SOTA SLAM system FAST-LIVO2 [10757429] is used to extract RGBD data and poses to create an instance-level open-vocabulary map. Building on an instance-level open-vocabulary map [werby23hovsg], we construct the Hierarchical Multi-modal Scene Graph (HMSG), which divides the environment into four levels-floor, room, view, and object. Each node is enriched with multi-modal features, including geometric attributes, semantic information, and topological connections.
During online interaction, the speech recognition module [gao2023funasr] converts user speech into text, which is then parsed by an LLM to extract the corresponding query. In the FSR pipeline, the query text is first grounded to candidate regions and objects through CLIP-based similarity matching and then refined via VLM-based visual verification to ensure robust and accurate navigation. By combining the HMSG representation with fast-to-slow reasoning (FSR), FSR-VLN enables efficient and interpretable navigation, particularly in long-range and cluttered real-world environments. Further details of this mechanism are provided in Section III-C. After obtaining the target object’s coordinates, the system performs path planning and whole-body control to reach the goal.
@@ III-B Hierarchical Multi-modal Scene Graph Construction
Following HOVSG [werby23hovsg], we first construct an open-vocabulary map at the instance-level. Objects and views are associated with rooms based on geometric overlap in top-down maps, resulting in a scene graph enriched with both geometric and semantic features.
@@ IV-A Experiment Setup
User Instructions. To evaluate the generalization of our system across diverse instructions, we follow the experimental setup of MobilityVLA [xu2024mobility] and crowd-source 87 user instructions in four categories: Reasoning-Free (RF) with explicit targets, Reasoning-Required (RR) requiring implicit goal inference, Small Objects (SO), which are challenging due to limited visual footprints, and Spatial Target (ST) designed to test long-range memory using room-type information. The instructions span 23, 18, 15, and 14 object categories, respectively.
@@
Baselines. We evaluated FSR-VLN against several SOTA ObjNav methods using different map representations: CLIP-based 3D voxel maps (OK-Robot [Liu_2024]), CLIP-based 3D scene graphs (HOVSG [werby23hovsg]), and image-based topological graphs (MobilityVLA [xu2024mobility]). OK-Robot performs object retrieval using CLIP-based OWL-ViT [minderer2022simpleopenvocabularyobjectdetection] features. HOVSG retrieves floors, rooms, and objects based on CLIP features, with its 3D scene graph structure supporting long-range retrieval. MobilityVLA retrieves goal images from all views; as it is not open-sourced, we implemented a version using GPT-4o, first selecting the top 50 candidate frames via CLIP similarity before inference due to GPT-4o’s context limitations. OK-Robot and MobilityVLA lack room-level information and cannot interpret Spatial Target (ST) instructions referencing a specific room. For fair comparison, we consider only the object query derived from LLM-based instruction understanding for these methods and HOVSG. Retrieval is also counted as successful if objects of the same type from other rooms are returned for ST instructions. In contrast, FSR-VLN leverages room-level spatial information, and retrieval is considered successful only when the target object is in the specified room. By integrating the HMSG map with navigation reasoning, FSR-VLN jointly predicts the goal view and object, and a retrieval is successful only if both are correctly identified.
In the HM3D-SEM dataset, as osmAG-LLM, we select HOV-SG and osmAG-LLM as our baseline because they share fundamental design principles with our approach[xie2025osmag].
@@ IV-B Benchmark Results
OK-Robot[jin2025openfusionopenvocabularyrealtimescene] CLIP based 3D Voxel Map 0.2s 0.609 0.609 0.609 0.609 0.609 0.609 0.632 0.632 0.632 0.632 0.632
HOVSG[werby23hovsg] CLIP based 3D Scene Graph 0.2s 0.517 0.517 0.573 0.586 0.596 0.596 0.770 0.816 0.828 0.828 0.828
MobilityVLA[xu2024mobility] Image based Topological Graph + VLM 30s 0.345 0.345 0.598 0.759 0.805 0.954 - - - - -